# An efficient Monte Carlo method for valid prior-free possibilistic statistical inference

### v2

@@ 3.1 Key insights
Of course, the normal distribution, ${\sf N}_{d}(m,V)$, assigns probability $1-\alpha$ to each $\alpha$-cut, so that must be the inner probabilistic approximation; in fact, this is a tautology since the Gaussian possibility measure above is defined as the outer possibilistic approximation of the Gaussian probability distribution (Martin and Williams, 2024). Following the guidance offered in the previous paragraph, one first finds that the marginal distribution $\mathsf{M}$ of $\pi(\Theta)$ is ${\sf Unif}(0,1)$ under $\Theta\sim{\sf N}_{d}(m,V)$. Second, since a Gaussian random vector—or any elliptically symmetric random vector for that matter (e.g., Hult and Lindskog, 2002, Theorem 3.1)—can be written as $\Theta=m+RV^{1/2}U$, where $V^{1/2}$ is the Cholesky factor of $V$, $(R,U)$ are independent, with $U$ a uniform random vector on the unit sphere in $\mathbb{R}^{d}$ and $R$ a non-negative random variable, the corresponding kernel $\mathsf{K}^{\beta}$ is the distribution of $m+\{F_{d}^{-1}(1-\beta)\}^{1/2}\,V^{1/2}\,U$, which is fully supported on the boundary $\partial C_{\beta}$ of $C_{\beta}$.
3.2 Probabilistic approximation to a possibilistic IM
Suppose I have at my disposal a collection of data-dependent probability distributions, $\{\mathsf{Q}_{z}^{\alpha}:\alpha\in[0,1]\}$, supported on $\mathbb{T}$, with the property
$\mathsf{Q}_{z}^{\alpha}\{C_{\alpha}(z)\}\geq 1-\alpha,\quad\alpha\in[0,1],$ (11)
where $C_{\alpha}(z)$ is the possibilistic IM’s $\alpha$-cut. Where specifically these $\mathsf{Q}_{z}^{\alpha}$’s might come from will be discussed in Section 3.3 below and in more detail in Appendix A.2. Note that the $\mathsf{Q}_{z}^{\alpha}$’s aren’t confidence distributions because no guarantees are offered concerning the $\mathsf{Q}_{z}^{\alpha}$-probability assigned to $C_{\beta}(z)$ for $\beta\neq\alpha$. Following the intuition provided by Theorem 1 and the subsequent example, I propose to stitch these $\alpha$-dependent probability distributions together into a single probability distribution, $\mathsf{Q}_{z}^{\star}$, using the mixture operation in (9). More specifically, I propose to take the marginal distribution $\mathsf{M}$ to be ${\sf Unif}(0,1)$ and the kernel/conditional distribution as follows: let $\Theta\sim\mathsf{Q}_{z}^{\alpha}$ and set the kernel $\mathsf{K}^{\alpha}\equiv\mathsf{K}_{z}^{\alpha}$, which now depends on data $z$, to be (a version of) the conditional distribution of $\Theta$, given $\Theta\in\partial C_{\alpha}(z)$, i.e.,
$\mathsf{K}_{z}^{\alpha}(H):=\mathsf{Q}_{z}^{\alpha}\{\Theta\in H\cap C_{\alpha}(z)\mid\Theta\in\partial C_{\alpha}(z)\}.$
Then my proposed probabilistic approximation of the possibilistic IM is based on stitching the $\alpha$-specific probabilities according to the following rule:
@@ 3.2 Probabilistic approximation to a possibilistic IM
Of course, evaluating features of or simulating from the distribution $\mathsf{Q}_{z}^{\star}$ in (12) could be challenging in general but, fortunately, this is relatively easy to do in the case of my recommended approach described in Section 3.3 below. At least intuitively, however, it’s clear how to proceed from (12) to evaluate the proposed probabilistic approximation of the IM: samples from $\mathsf{Q}_{z}^{\star}$ can be obtained by first sampling $A\sim{\sf Unif}(0,1)$ and then by sampling $\Theta$ from the conditional distribution $\mathsf{Q}_{z}^{A}$ that’s concentrated on the boundary $\partial C_{A}(z)$. If there’s a specific feature of $\mathsf{Q}_{z}^{\star}$ to be evaluated, then it may be advantageous to employ a Rao–Blackwellization strategy to reduce the Monte Carlo variance. That is, suppose the goal is to evaluate the expected value, say, $\mathsf{Q}_{z}^{\star}h$, of some real-valued function $h$ defined on $\mathbb{T}$, such as $h=1_{H}$ for a relevant hypothesis $H$. Then first get $\alpha\mapsto\mathsf{K}_{z}^{\alpha}h$, the expected value of $h(\Theta)$ relative to the conditional distribution $\mathsf{K}_{z}^{\alpha}$ using Monte Carlo, and then integrate over $\alpha$ numerically, e.g., using R’s integrate function.
Keep in mind that the possibilistic IM is the target, and even the “best” probabilistic approximation can be far off from the target in its answers to certain questions. In general, all that’s guaranteed is $\mathsf{Q}_{z}^{\star}(H)\leq\overline{\Pi}_{z}(H)$, and equality holds effectively only when $H$ is the complement of an $\alpha$-cut. Similar statements can be made for expectations of general functions $h$, but it’s a bit more complicated: basically, the probabilistic $\mathsf{Q}_{z}^{\star}$-expectations agree with the possibilistic $\overline{\Pi}_{z}$-upper expectations when $h$ has level sets that match the IM’s collection of $\alpha$-cuts. Section 3.5 below offers an approach to overcome the inherent limitations of any probabilistic approximation of the possibilistic IM.
3.3 Implementation details
Cella and Martin, (2024) proposed introducing a parametric family, say, $\{\mathsf{R}_{z}^{\xi}:\xi\in\Xi\}$ of probability distributions on $\mathbb{T}$ and, akin to variational Bayes analysis (e.g., Blei et al., 2017), choosing a particular value $\xi=\xi(z,\alpha)$ of the parameter—depending on data $z$ and level $\alpha$—such that $\mathsf{Q}_{z}^{\alpha}:=\mathsf{R}_{z}^{\xi(z,\alpha)}$ assigns at least $1-\alpha$ probability to the $\alpha$-cut $C_{\alpha}(z)$, as required in (11). Intuitively, if $\xi$ controls the spread of $\mathsf{R}_{z}^{\xi}$, and if this distribution is properly centered and oriented, then there exists a value $\xi(z,\alpha)$ such that the $\mathsf{Q}_{z}^{\alpha}$ defined in the previous sentence has the stated property, at least approximately. As the reader might expect, some problem-specific considerations are required to efficiently implement this strategy, but it can be done accurately in a wide range of applications as demonstrated next and in the reference cited above.
For the relative likelihood-based possibilistic IM described in Section 2, under the usual regularity conditions, Martin and Williams, (2024) established an asymptotic Gaussianity result that says, if the sample size is large, then
@@ 3.3 Implementation details
where $\hat{\theta}_{z}$ is the maximum likelihood estimator, $J_{z}$ is the $d\times d$ observed Fisher information matrix, and, again, $F_{d}$ is the ${\sf ChiSq}(d)$ distribution function. This Gaussian approximation also holds in alternative parametrizations and, in fact, it might be more accurate in certain parametrizations than in others. For example, if there are non-negativity constraints on $\theta$, then the Gaussian approximation might be more appropriate/accurate when applied on the $\log\theta$ scale than on the $\theta$ scale. My general presentation below works on the $\theta$ scale, but I will apply these transformations freely in the examples that follow.
Since the best probabilistic approximation of the Gaussian possibility measure is the familiar Gaussian distribution, a reasonable choice of the variational family is
$\mathsf{R}_{z}^{\xi}={\sf N}_{d}\bigl{(}\hat{\theta}_{z},\,J_{z}^{-1}(\xi)\bigr{)},\quad\xi\in\Xi=(0,\infty)^{d}$ (14)
where, if $J_{z}=E\Lambda E^{\top}$ is the spectral decomposition of $J_{z}$, then
$J_{z}(\xi)=E\,\text{diag}(\xi^{-1})\,\Lambda\,\text{diag}(\xi^{-1})\,E^{\top},$
for $\xi^{-1}$ the entry-wise reciprocal of $\xi$, and $\text{diag}(\cdot)$ the operator that takes its vector argument to a diagonal matrix with that vector on the diagonal. Now it should be clear what was meant when I said above that “$\xi$ controls the spread of $\mathsf{R}_{z}^{\xi}$.” With this family of candidate probabilistic approximations to the possibilistic IM, Cella and Martin, (2024) proposed a stochastic approximation algorithm (see Appendix A.2 below) that, for a given $\alpha$, selects $\xi=\xi(z,\alpha)$ such that
$\mathsf{R}_{z}^{\xi(z,\alpha)}\{C_{\alpha}(z)\}\geq 1-\alpha,\quad\text{at least approximately},$
where $\mathsf{R}_{z}^{\xi(z,\alpha)}$ is the Gaussian distribution in (14) with $\xi=\xi(z,\alpha)$ plugged into the variance and $C_{\alpha}(z)$ is still the original IM’s $\alpha$-cut. More specifically, if the variational family’s $\alpha$-cuts are defined as
$C_{\alpha}^{\xi}(z)$ =$=\{\theta:1-F_{d}\bigl{(}(\theta-\hat{\theta}_{z})^{\top}J_{z}(\xi)(\theta-\hat{\theta}_{z})\bigr{)}\geq\alpha\}$
=$=\{\theta:(\theta-\hat{\theta}_{z})^{\top}J_{z}(\xi)(\theta-\hat{\theta}_{z})\leq F_{d}^{-1}(1-\alpha)\},$ (15)
then Cella and Martin’s algorithm returns $\xi=\xi(z,\alpha)$ such that
$K_{\alpha}^{\xi(z,\alpha)}(z)\supseteq C_{\alpha}(z).$
They basically stopped here, suggesting $\mathsf{Q}_{z}^{\alpha}=\mathsf{R}_{z}^{\xi(z,\alpha)}$ as a limited-but-simple probabilistic approximation of the possibilistic IM’s output $\overline{\Pi}_{z}$ with the property of being a confidence distribution at least for the specified confidence level $\alpha$.
To move beyond the insights in Cella and Martin, the key two-fold observation here is as follows: first, $C_{\alpha}^{\xi}(z)$ has a simple expression and geometric form (ellipsoid) and, second, it is easy to condition the Gaussian distribution $\mathsf{Q}_{z}^{\alpha}$ to the boundary $\partial C_{\alpha}^{\xi(z,\alpha)}(z)$—this is exactly what was shown in the Gaussian example at the end of Section 3.1. Of course, bigger $\alpha$-cuts imply more conservative inference, but empirical evidence in Cella and Martin, (2024) and Section 4 below suggests that the approximation is quite accurate, i.e., the two $\alpha$-cuts in the above display are nearly the same. Just a slight loss of statistical efficiency is a small price to pay for replacing the impossible task of conditioning on $\partial C_{\alpha}(z)$ with the easy task of conditioning on $\partial C_{\alpha}^{\xi(z,\alpha)}(z)$.
To summarize, I propose to implement the probabilistic approximation $\mathsf{Q}_{z}^{\star}$ in (12) by iterating the following two steps:
@@
    Given $A$, evaluate $\xi(z,A)$, and then sample $\Theta$ just as described in the Gaussian example at the end of Section 3.1, i.e.,
@@
Of course, both sampling steps above are trivial; the only expensive operation—“evaluate $\xi(z,A)$”—is almost hidden in the description. The evaluation of $\xi(z,A)$ involves an $A$-specific run of the stochastic approximation algorithm summarized in Appendix A.2 which, fortunately, typically only involves a few updates and a few Monte Carlo evaluations of the IM contour $\pi_{z}$ per update. So, this is fast and easy to compute, and it scales linearly in dimension. Moreover, it’s easy to parallelize the evaluation of $\xi(z,A)$ over multiple $A$’s, dramatically reducing computation time.
A practical adjustment to the above procedure, which allows for larger sample sizes from $\mathsf{Q}_{z}^{\star}$ with even less computation time, is as follows. Start with a fixed grid $\mathcal{A}$ of, say 100 $\alpha$ values; in my examples that follow, I use an equally-spaced grid between 0.001 and 0.999. Then evaluate $\xi(z,\alpha)$ for all $\alpha\in\mathcal{A}$, which sets a fixed limit on the amount of computational investment required. Now, for each $A\sim{\sf Unif}(0,1)$ drawn in Step 1 above, find the two $\alpha$ values in $\mathcal{A}$ that sandwich $A$ and get $\xi(z,A)$ via linear interpolation, and then do Step 2 as described above. Once the 100 values of $\xi(z,\cdot)$ are obtained, which takes roughly 5 seconds in my examples below, the samples from $\mathsf{Q}_{z}^{\star}$ are virtually free.
Finally, the procedure proposed here has certain aspects in common with the calibrated bootstrap algorithm developed in Jiang et al., (2023); in fact, the results in this paper and in Cella and Martin, (2024) were inspired by the calibrated bootstrap efforts. On the one hand, the calibrated bootstrap proposal is more general than that here because it doesn’t require specification of a parametric/Gaussian family of approximate distributions—it uses bootstrap to learn a full distribution nonparametrically. On the other hand, learning the full distribution requires computational effort and, furthermore, that full distribution is not well-suited for the condition-to-the-boundary step. Consequently, the calibrated bootstrap will spend considerable computational resources to evaluate the contour $\pi_{z}$ at points which will ultimately be discarded from the sample that describes $\mathsf{Q}_{z}^{\star}$.
3.4 Asymptotic analysis
Here I offer a brief large-sample analysis to justify the claimed overall accuracy of the proposed probabilistic approximation. For the sake of brevity, and without loss of persuasiveness, I give only a heuristic argument building on the asymptotic convergence results rigorously demonstrated in Martin and Williams, (2024).
As mentioned above, Martin and Williams, (2024) show that, under the regularity conditions sufficient for asymptotic normality and efficiency of the maximum likelihood estimator, the relative likelihood-based possibilistic IM enjoys a large-sample Gaussianity property, akin to the classical Bernstein–von Mises theorem fundamental to Bayesian analysis; see (13). The uniform mode of their convergence result implies, first, that the IM’s $\alpha$-cuts, $C_{\alpha}(z)$, will merge with the $\alpha$-cuts, $C_{\alpha}^{1}(z)$, of the variational approximation, with $\xi=1$, as the size of the sample $z$ increases to $\infty$. Second, since $\xi(z,\alpha)$ is designed to make the variational family’s $\alpha$-cuts agree with the original IM’s, one fully expects that the components of $\xi(z,\alpha)$ are converging to 1 as the sample size increases.
Putting everything together, the original IM’s $\alpha$-cuts merge with the corresponding Gaussian $\alpha$-cuts asymptotically, and the components of the variational family index $\xi(z,\alpha)$ identified by Cella and Martin’s algorithm are converging to 1. Therefore, the symmetric difference $C_{\alpha}(z)\,\triangle\,C_{\alpha}^{\xi(z,\alpha)}(z)$ is converging to $\varnothing$, which implies that the proposed sampling algorithm exactly recovers the IM’s limiting inner probabilistic approximation, which is the Gaussian distribution just like in the example in Section 3.1.

### v3

@@ 3.1 Probabilistic approximation to a possibilistic IM
Of course, the normal distribution, ${\sf N}_{d}(m,V)$, assigns probability $1-\alpha$ to each $\alpha$-cut, so that must be the inner probabilistic approximation; in fact, this is a tautology since the Gaussian possibility measure is defined as the outer possibilistic approximation of the Gaussian probability distribution (Martin and Williams, 2025). For a direct connection to Corollary 1, one first finds that the marginal distribution $\mathsf{M}$ of $\pi(\Theta)$ is ${\sf Unif}(0,1)$ under $\Theta\sim{\sf N}_{d}(m,V)$. Second, since a Gaussian random vector—or any elliptically symmetric random vector for that matter (e.g., Hult and Lindskog, 2002, Theorem 3.1)—can be written as $\Theta=m+RV^{1/2}U$, where $V^{1/2}$ is the Cholesky factor of $V$, $(R,U)$ are independent, with $U$ a uniform random vector on the unit sphere in $\mathbb{R}^{d}$ and $R$ a non-negative random variable, the corresponding kernel $\mathsf{K}^{\beta}$ is the distribution of $m+\{F_{d}^{-1}(1-\beta)\}^{1/2}\,V^{1/2}\,U$, which is fully supported on the boundary $\partial C_{\beta}$ of $C_{\beta}$.
A direct implementation of the inner probabilistic approximation of the possibilistic IM based on the characterization in Corollary 1 is given in Appendix C. That direct approximation would generally be more expensive computationally compared to using a second-layer approximation as described in Sections 3.23.3. No claims are made that this second-layer approximation is universal but, for the regular parametric models under consideration here, the results shown below based on this simpler approach are strikingly accurate, so the computationally cheaper solution is preferred.
3.2 Approximating the probabilistic approximation
To set the scene for a second-layer approximation, suppose that, in addition to the possibilistic IM output $\pi_{z}$ and $\Pi_{z}$, I have another data-dependent possibility contour $\pi_{z}^{\text{out}}$ and corresponding possibility measure $\Pi_{z}^{\text{out}}$. These two sets of possibilistic output are related in the following critical way: $\pi_{z}(\theta)\leq\pi_{z}^{\text{out}}(\theta)$ for all $\theta\in\mathbb{T}$. The construction of such a $\pi_{z}^{\text{out}}$ will be discussed in Section 3.3. Ideally, $\pi_{z}$ and $\pi_{z}^{\text{out}}$ would be close, subject to the dominance constraint, so I’ll interpret $\pi_{z}^{\text{out}}$ as an outer approximation of $\pi_{z}$. The dominance property “$\pi_{z}\leq\pi_{z}^{\text{out}}$” implies that $\Pi_{z}$ is more specific than $\Pi_{z}^{\text{out}}$, i.e., $\Pi_{z}$ makes stronger claims than does $\Pi_{z}^{\text{out}}$. The interpretation of specificity might be easiest to understand through the relation induced between the credal sets:
$\pi_{z}\leq\pi_{z}^{\text{out}}\implies\mathscr{C}(\Pi_{z})\subseteq\mathscr{C}(\Pi_{z}^{\text{out}}).$
In words, it takes fewer probability distributions to describe $\Pi_{z}$ than to describe $\Pi_{z}^{\text{out}}$; consequently, there’s less (epistemic) uncertainty in $\Pi_{z}$ than in $\Pi_{z}^{\text{out}}$ and, hence, the former makes stronger claims than the latter. It suffices to interpret $\Pi_{z}^{\text{out}}$ as an IM that’s (slightly) more conservative than $\Pi_{z}$. Importantly, since the original IM with contour $\pi_{z}$ is valid in the sense of (3), then so is the IM with contour $\pi_{z}^{\text{out}}$. For example, the $\alpha$-cuts $C_{\alpha}^{\text{out}}(z)=\{\theta:\pi_{z}^{\text{out}}(\theta)\geq\alpha\}$ of $\pi_{z}^{\text{out}}$ are confidence sets, just like $C_{\alpha}(z)$. $\Pi_{z}^{\text{out}}$ is appealing only if its computational benefits outweigh the loss of specificity/efficiency compared to $\Pi_{z}$. This is often the case, as the numerical results below show.
From here, my proposal is straightforward: approximate $\mathsf{Q}_{z}^{\text{inn}}$ by $\mathsf{Q}_{z}^{\star}$, where the latter is defined as an inner probabilistic approximation of the outer possibilistic approximation $\Pi_{z}^{\text{out}}$. The relationship between $\mathsf{Q}_{z}^{\text{inn}}$ and $\mathsf{Q}_{z}^{\star}$ can be described as follows:
$\mathsf{Q}_{z}^{\text{inn}}\in\mathscr{C}(\Pi_{z})\subseteq\mathscr{C}(\Pi_{z}^{\text{out}})\ni\mathsf{Q}_{z}^{\star}.$
From this it’s at least intuitively clear that, if $\pi_{z}^{\text{out}}\approx\pi_{z}$ (subject to the dominance constraint), then $\mathsf{Q}_{z}^{\star}$ and $\mathsf{Q}_{z}^{\text{inn}}$ should be similarly close. Note, however, that $\mathsf{Q}_{z}^{\star}\not\in\mathscr{C}(\Pi_{z})$ in general. That $\mathsf{Q}_{z}^{\star}$ might fail to be a member of $\mathscr{C}(\Pi_{z})$ is not an issue. What matters is that $\mathsf{Q}_{z}^{\star}$ can maintain the reliability of $\Pi_{z}$ (or $\mathsf{Q}_{z}^{\text{inn}}$) while being easier to compute. In fact, since all the members of $\mathscr{C}(\Pi_{z})$ are, by definition, at least as tightly concentrated as $\mathsf{Q}_{z}^{\text{inn}}$, an arguably safer strategy is to expand the set $\mathscr{C}(\Pi_{z})$ of candidate approximations a bit to create a better opportunity of finding one with the “right” concentration.
Since $\mathsf{Q}_{z}^{\star}$ is itself an inner probabilistic approximation, a concrete expression for it, along with a strategy for numerically approximating it via Monte Carlo follows from the characterization result in Corollary 1. Indeed, $\mathsf{Q}_{z}^{\star}$ can be written as
@@ 3.2 Approximating the probabilistic approximation
where $\mathsf{K}_{z}^{\alpha}$ is a distribution supported on $\partial C_{\alpha}^{\text{out}}(z)$. Whether $\mathsf{Q}_{z}^{\star}$ in (10) is an accurate approximation of and easier to compute than $\mathsf{Q}_{z}^{\text{inn}}$ depends on the choice of $\pi_{z}^{\text{out}}$. Section 3.3 presents a simple construction of $\pi_{z}^{\text{out}}$ that achieves both of these desiderata.
3.3 Implementing the second-layer approximation
Here I construct the possibility contour $\pi_{z}^{\text{out}}$ that dominates $\pi_{z}$ in the sense described in Section 3.2. Since the contour $\pi_{z}^{\text{out}}$ is determined by its $\alpha$-cuts $\{C_{\alpha}^{\text{out}}(z):\alpha\in[0,1]\}$, the presentation that follows focuses on the latter.
For the relative likelihood-based possibilistic IM defined in Section 2, under the usual regularity conditions, Martin and Williams, (2025) established an asymptotic Gaussianity result that says, if the sample size is large, then
@@ 3.3 Implementing the second-layer approximation
where $\hat{\theta}_{z}$ is the maximum likelihood estimator, $J_{z}$ is the $d\times d$ observed Fisher information matrix, and, again, $F_{d}$ is the ${\sf ChiSq}(d)$ distribution function. Consequently, the $\alpha$-cuts $C_{\alpha}(z)$ of $\pi_{z}$ are approximately ellipsoidal. This suggests a basic elliptical form for $C_{\alpha}^{\text{out}}(z)$ and hence for $\pi_{z}^{\text{out}}$, which is asymptotically correct/optimal, and is shown to be quite accurate even for small samples. As I’ll explain below, assuming that $C_{\alpha}(z)$ can be roughly approximated by an ellipsoid is not equivalent to assuming Gaussianity.
Note that the approximation (11) also holds for different parametrizations. In fact, it might be more accurate in one parametrization than in another, e.g., non-negativity constraints on $\theta$ suggest applying a Gaussian approximation on the $\log\theta$-scale. What’s described below on the $\theta$-scale can also be applied following a reparametrization, and I freely reparametrize as appropriate in my numerical examples below.
For constructing $C_{\alpha}^{\text{out}}(z)$, I leverage an idea advanced by Cella and Martin, (2025). Below I describe their idea—different from how they described it—and the relevant technical details are given in Appendix D. Define an eigenvalue-scaled version $J_{z}(\sigma)$ of the observed information $J_{z}$ as follows: if $J_{z}=E\Lambda E^{\top}$ is the spectral decomposition, then
$J_{z}(\sigma)=E\,\text{diag}(\sigma^{-1})\,\Lambda\,\text{diag}(\sigma^{-1})\,E^{\top},$
where $\sigma^{-1}$ is the entry-wise reciprocal of the vector $\sigma\in(0,\infty)^{d}$, and $\text{diag}(\cdot)$ is the operator that takes a vector argument to a diagonal matrix with that vector on the diagonal. Then $\sigma$ acts like a vector of scaling parameters that can stretch or contract the level sets of the quadratic form $\theta\mapsto(\theta-\hat{\theta}_{z})^{\top}J_{z}(\sigma)^{-1}(\theta-\hat{\theta}_{z})$ in different directions. Next, define the $\sigma$-specific $\alpha$-cut for a Gaussian ${\sf N}_{d}(\hat{\theta}_{z},J_{z}(\sigma)^{-1})$ possibility contour
$C_{\alpha}^{\sigma}(z)$ =$=\{\theta:1-F_{d}\bigl{(}(\theta-\hat{\theta}_{z})^{\top}J_{z}(\sigma)(\theta-\hat{\theta}_{z})\bigr{)}\geq\alpha\}$
=$=\{\theta:(\theta-\hat{\theta}_{z})^{\top}J_{z}(\sigma)(\theta-\hat{\theta}_{z})\leq F_{d}^{-1}(1-\alpha)\},$ (12)
Since $C_{\alpha}(z)$ is, for each $\alpha$ and $z$, approximately elliptical with center $\hat{\theta}_{z}$ and shape roughly that of a $J_{z}$-quadratic form, one expects that there exists a scaling vector $\sigma=\sigma(z,\alpha)$ such that, at least approximately,
$C_{\alpha}^{\sigma(z,\alpha)}(z)\supseteq C_{\alpha}(z),\quad\alpha\in[0,1].$
A cartoon illustration of this bounding is shown in Figure 1, and the algorithm proposed in Cella and Martin, (2025) for finding such a $\sigma(z,\alpha)$ is given in Appendix D. There are two important points worth mentioning here. First, the algorithm developed in Cella and Martin, (2025) is efficient in the sense that evaluating $\sigma(z,\alpha)$ has complexity is $O(d)$ compared to $O(e^{d})$ for the naive approach in (6). Second, note that it’s not a single scaling factor $\sigma$ that achieves the containment in the above display for each $\alpha$ as would be the case if $\pi_{z}$ were exactly Gaussian; instead, it’s an $\alpha$-specific adjustment which, at least intuitively, suggests more flexible “scale mixture of normals” form.
[…]
Figure 1: Approximating the IM contour’s $\alpha$-cut $C_{\alpha}(z)$ by an ellipse $C_{\alpha}^{\sigma}(z)$ with a “good” choice of $\sigma$ (solid) and with a “bad” choice of $\sigma$ (dashed).
Putting everything together, my specific proposal is to define the $C_{\alpha}^{\text{out}}(z):=C_{\alpha}^{\sigma(z,\alpha)}(z)$, which, in turn, leads to the probabilistic approximation $\mathsf{Q}_{z}^{\star}$ as in (10). And the mixture form (10) immediately suggests a Monte Carlo method for evaluating various $\mathsf{Q}_{z}^{\star}$-probabilities; see Section 3.4. The whole procedure is succinctly described in Algorithm 1. A brief asymptotic analysis justifying the accuracy of the proposed approximation is presented in Remark 2 of Appendix E. A few conceptual points are given next.
  • Of course, the two sampling steps in Algorithm 1 are trivial. The only non-trivial step, namely “evaluate $\sigma(z,\alpha)$,” is easy to miss. This involves an $\alpha$-specific run of the algorithm summarized in Appendix D which, fortunately, typically only involves a few updates and a few Monte Carlo evaluations of the IM contour $\pi_{z}$ per update. So, this is easy to compute and has complexity scaling linearly in dimension $d$. Moreover, it’s easy to parallelize the evaluation of $\sigma(z,\alpha)$ over multiple $\alpha$’s, dramatically reducing computation time; see, also, Remark 1 in Appendix E.
  • The benefit to using the elliptical approximations should now be apparent, since sampling from $\mathsf{Q}_{z}^{\star}$ is effectively no different than sampling from a Gaussian as in the illustration towards the end of Section 3.1. But remember that $\mathsf{Q}_{z}^{\star}$ is generally not Gaussian—the $\alpha$-adaptivity gives it more flexibility.
  • Despite the aforementioned adaptivity, forcing $\partial C_{\alpha}^{\text{out}}(z)$ to be elliptically-shaped potentially implies some inefficiency in small samples, as Figure 1 reveals. Theoretically and empirically, this restriction is rather mild, but there may be applications in which an elliptical shape might not be appropriate. Fortunately, the elliptical structure isn’t necessary to deploy the solution proposed here; see Appendix C.
Independent draws from $\mathsf{Q}_{z}^{\star}$ in (10) are obtained by repeating the following steps:
@@
    Given $A=\alpha$, evaluate $\sigma(z,\alpha)$, and then sample $\Theta$ just as described in the Gaussian example at the end of Section 3.1, i.e.,
@@
Algorithm 1 Sampling from the probabilistic approximation $\mathsf{Q}_{z}^{\star}$ of $\Pi_{z}$.
Two final remarks to put the contribution into perspective will close out this subsection. First, Cella and Martin’s goal was the same as that in the present paper. What they offered as their final, albeit incomplete solution was a simple, $\alpha$-specific Gaussian probabilistic approximation, namely, $\mathsf{Q}_{z}^{\text{\sc cm},\alpha}={\sf N}_{d}(\hat{\theta}_{z},J_{z}(\sigma(z,\alpha))^{-1})$, which is then transformed into a possibility measure as described in Section 3.4 below. Their approximation is only designed to capture a certain aspect of the original IM, in particular, its $\alpha$-cut for the one specified $\alpha$, so it need not be (although often is) a good approximation of $\pi_{z}$ overall. The goal is obviously a good overall approximation to $\pi_{z}$ and Cella and Martin, (2025, Sec. 6) acknowledge this limitation: “finding a way to stitch together these $\alpha$-specific IM approximations is an important open problem.” The present paper’s contribution is filling this significant gap in the IM computational developments.
Second, my proposed solution has certain aspects in common with the calibrated bootstrap developed in Jiang et al., (2024). In fact, my solution was actually inspired by the calibrated bootstrap. Specialized points are needed to make this comparison so, for the sake of brevity here, I defer these details to Remark 3 in Appendix E.